OpenStudy (anonymous):

what is f(x) when the lim x approaches negative infinity equals 2?

OpenStudy (anonymous):

could be lots of different things

OpenStudy (anonymous):

I just need one possible answer

OpenStudy (bahrom7893):

well a possible answer could be 2x/x

OpenStudy (bahrom7893):

or (2x+1)/x

OpenStudy (bahrom7893):

or (2x-pi)/(x-(4/3))

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

how about \(2+\frac{\cos(x)}{x}\)

OpenStudy (bahrom7893):

That too lol

OpenStudy (anonymous):

Okay like for this problem I have to have a y=f(x) that fits all of these different limits but it can be more than one function like with asymptotes and stuff.. can you help me?!